In quebec, we have to pay 2 taxes.Federale taxe:6%. Qucebec tave:7,5%.Why are these 2 combined taxes give a pecentage of 13,95% and not 13,5%?????????? PLESEEEEEEEEEEEEEEEEEEEEEE EXPLAINNNNNNNNNNN !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
Answer by ankor@dixie-net.com(22740)

In Quebec, we have to pay 2 taxes. Federale taxe:6%. Quebec taxe:7,5%. Why are these 2 combined taxes give a percentage of 13,95% and not 13,5%
;
Perhaps they apply the Quebec tax after the Federal tax has be added:
:
for example use a $100 purchase:
Obviously, after a 6% Federal tax was added, it would be a total of $106
Then they apply the provincial tax to that. .075 * 106 = 7.95.
That is added so 106 + 7.95 = $113.95 is the total
13.95 total taxes, Find the tax rate 13.95/100 * 100 = 13.95%
:
Did this make sense? Hope so.